# EEG-SCMM: Soft Contrastive Masked Modeling for Cross-Corpus EEG-Based Emotion Recognition

### v1

@@ Appendix D Masking Strategy
This paper introduces a novel hybrid masking strategy to generate diverse masked samples by considering both channel and feature relationships. To compare our approach with traditional masking strategies, we explore three different masking rules: random masking, channel masking, and hybrid masking. Figure 3 illustrates the difference between these three masking strategies.
@@
(1) Random Masking: Generate masks using a binomial distribution to randomly mask samples in the feature dimension, setting the values of the masked features to zero.
@@
(2) Channel Masking: Generate masks using a binomial distribution to randomly mask samples in the channel dimension, setting the values of all features within the masked channels to zero.
@@
(3) Hybrid Masking: Generate a probability matrix using a uniform distribution that proportionally mixes masks generated from random masking and channel masking.
Our proposed hybrid masking strategy is highly flexible and can be extended to various datasets by integrating multiple masking strategies in different ratios, which is exceptionally suitable for data with rich semantic information. This approach effectively generates more diverse masked samples, encouraging the model to comprehensively capture the inherent relationships of the data.
Appendix E Hyperparameter Analysis
We conduct comprehensive experiments to verify the hyperparameter sensitivity of SCMM on the SEED and SEED-IV 3-category datasets. The hyperparameters examined include the masking ratio $r$, probability threshold $\mu$, metric function $Dist(\cdot,\cdot)$, upper bound $\alpha$, sharpness $\tau_{w}$, and temperature $\tau_{c}$. The complete experimental results are presented in the following sections.
E.1 Masking Ratio
Table 9 shows the model performance in the SEED-IV3 $\rightarrow$ SEED3 and SEED3 $\rightarrow$ SEED-IV3 experiments when using different masking ratios (0.1, 0.25, 0.5, and 0.75). Experimental results demonstrate that the proposed SCMM consistently achieves the best performance when the masking ratio is set to $r=0.5$.
Ratio ACC / STD (%)
SEED-IV3 $\rightarrow$ SEED3 SEED3 $\rightarrow$ SEED-IV3
@@ E.1 Masking Ratio
Table 9: Hyperparameter sensitivity analysis of masking ratio $r$ on SEED and SEED-IV.
E.2 Probability Threshold
Table 10 presents the model performance using different probability thresholds (ranging from 0 to 1) for hybrid masking. Here, $\mu=0$ refers to the exclusive use of the random masking strategy, while $\mu=1$ indicates the exclusive use of the channel masking strategy. Experimental results show that SCMM achieves the best performance in the SEED-IV3 $\rightarrow$ SEED3 and SEED3 $\rightarrow$ SEED-IV3 experiments when the probability threshold is set to $\mu=0.1$ (i.e., the ratio of random masking and channel masking is 9:1).
Threshold ACC / STD (%)
SEED-IV3 $\rightarrow$ SEED3 SEED3 $\rightarrow$ SEED-IV3
@@ E.2 Probability Threshold
Table 10: Hyperparameter sensitivity analysis of threshold $\mu$ on SEED and SEED-IV.
E.3 Metric Function
We evaluate the model performance of SCMM using different metric functions in soft assignments $w(\cdot,\cdot)$. Experimental results in Table 11 indicate that SCMM performs best on the SEED and SEED-IV 3-category datasets when cosine similarity is used as the metric function.
Metrics ACC / STD (%)
SEED-IV3 $\rightarrow$ SEED3 SEED3 $\rightarrow$ SEED-IV3
@@ E.3 Metric Function
Table 11: Hyperparameter sensitivity analysis of metric function $Dist(\cdot,\cdot)$ on SEED and SEED-IV.
E.4 Upper Bound
We assess the impact of different upper bounds on calculating soft assignments $w(\cdot,\cdot)$. To restrict the soft assignments to the range of 0 to 1, we explore the model performance of SCMM with upper bounds of 0.25, 0.5, 0.75, and 1. Experimental results in Table 12 illustrate that the proposed SCMM achieves the best performance in the SEED-IV3 $\rightarrow$ SEED3 experiment when the upper bound is set to $\alpha=0.5$.
Upper Bound ACC / STD (%)
SEED-IV3 $\rightarrow$ SEED3
0.25 91.21 / 08.54
0.5 91.61 / 07.56
0.75 91.12 / 08.25
1 90.50 / 08.00
Table 12: Hyperparameter sensitivity analysis of upper bound $\alpha$ in the SEED-IV3 $\rightarrow$ SEED3 experiment.
E.5 Sharpness
Table 13 explore the impact of different sharpness parameters in the SEED-IV3 $\rightarrow$ SEED3 experiment, as shown in Table 13. Experimental results indicate that our model achieves the best performance with a sharpness parameter of $\tau_{w}=0.05$.
Sharpness ACC / STD (%)
SEED-IV3 $\rightarrow$ SEED3
0.01 90.93 / 09.13
0.05 91.61 / 07.56
0.1 91.08 / 09.00
0.2 90.59 / 07.40
0.5 90.63 / 09.16
1 90.09 / 08.62
2 89.58 / 09.26
Table 13: Hyperparameter sensitivity analysis of sharpness $\tau_{w}$ in the SEED-IV3 $\rightarrow$ SEED3 experiment.
E.6 Temperature
We conduct comparison experiments on the SEED and SEED-IV 3-category datasets to explore the impact of temperature parameter on model performance. Experimental results in Table 14 show that SCMM consistently achieves the best performance in both experiments when the temperature parameter is set to $\tau_{c}=0.5$.
Temperature ACC / STD (%)
SEED-IV3 $\rightarrow$ SEED3 SEED3 $\rightarrow$ SEED-IV3
@@ E.6 Temperature
Table 14: Hyperparameter sensitivity analysis of temperature $\tau_{c}$ on SEED and SEED-IV.
F Model Performance and Computational Complexity Analysis of SCMM
Metrics 1% of labeled data
ACC / STD (%) 74.98 / 17.00 71.35 / 16.29 73.20 / 16.09 60.59 / 23.72
Time Cost (s) 232.35 97.11 232.35 130.47
5% of labeled data
ACC / STD (%) 95.23 / 06.98 87.20 / 16.12 90.51 / 12.43 86.01 / 09.76
Time Cost (s) 240.48 100.23 240.48 132.90
10% of labeled data
ACC / STD (%) 99.86 / 00.31 90.04 / 16.94 99.72 / 00.80 94.43 / 05.60
Time Cost (s) 250.81 102.03 250.81 135.67
20% of labeled data
ACC / STD (%) 100.0 / 00.00 94.16 / 08.52 100.0 / 00.00 98.91 / 02.62
Time Cost (s) 267.21 105.48 267.21 139.31
Model Parameters: 1.52 M
Table 15: Model performance and computational complexity of SCMM when fine-tuning with limited labeled data (1%, 5%, 10%, and 20%) on SEED and SEED-IV.
To investigate the trade-off between model performance and computational complexity in few-shot scenarios, we assess the classification accuracy, time cost, and the number of trainable parameters of SCMM when fine-tuning with limited labeled data. Experimental results in Table 15 show that the model performance of SCMM significantly improves as the number of fine-tuning labeled samples increases. Meanwhile, the inference time cost remains low due to our lightweight design. Additionally, our model achieves superior classification performance in cross-corpus EEG-based emotion recognition tasks with very few parameters. In summary, the proposed SCMM effectively balances model performance and computational complexity.
G Visualization
G.1 Intra- and Inter-Class Similarities
Figure 7: Intra- and inter-class cosine similarities of embeddings learned by hard CL and soft CL. (a) is the average intra-class cosine similarity, and (b) is the average inter-class cosine similarity. Best viewed in color.
To assess the quality of the embeddings learned by SCMM, we randomly select one subject from the SEED dataset, and calculate both the average intra- and inter-class cosine similarities between the learned embeddings of all test samples, as shown in Fig. 7. It is evident that the proposed SCMM produces embeddings with higher intra-class similarity compared to traditional hard contrastive learning. In addition, the average inter-class similarity of the embeddings learned by SCMM is significantly lower than that of hard CL. In summary, visualization results confirm that the soft contrastive learning strategy designed in SCMM effectively clusters samples within the same category while distinctly separating samples from different categories, thus enhancing the model’s discriminative capabilities.
G.2 Reconstruction Quality
Figure 8: Comparison of reconstruction quality. We visualize the reconstruction results to compare the single-sample reconstruction paradigm (MAE) with the aggregate reconstruction paradigm (SCMM). This case is from the DEAP dataset. Best viewed in color.
To verify the effectiveness of the aggregator designed in SCMM, we compare the reconstruction quality of the single-sample reconstruction paradigm (MAE) with the aggregate reconstruction paradigm (SCMM) on the DEAP dataset. For clarity, we flatten both the original input sample and the reconstructed sample into one-dimensional vectors with dimensions $C\times F$ (channels $\times$ features). The results depicted in Fig. 8 illustrate that our model achieves lower reconstruction loss (MSE) and better sample reconstruction.

### v2

@@ Appendix D Masking Strategy
This paper introduces a novel hybrid masking strategy to generate diverse masked samples by considering both channel and feature relationships. To compare our approach with traditional masking strategies, we explore three different masking rules: random masking, channel masking, and hybrid masking. Figure 10 presents examples of generated masked samples using three strategies.
@@
  • Random Masking: Generate masks using a binomial distribution to randomly mask samples along the feature dimension, setting the values of masked features to 0.
@@
  • Channel Masking: Generate masks using a binomial distribution to randomly mask samples along the channel dimension, setting the values of all features within the masked channels to 0.
@@
  • Hybrid Masking: Generate a probability matrix using a uniform distribution that proportionally mixes masks generated by random masking and channel masking.
To assess the impact of various masking strategies, we conduct comparative experiments on the SEED and SEED-IV 3-category datasets using four strategies: random, channel, parallel, and hybrid masking. Specifically, the parallel masking strategy indicates that each sample is augmented randomly using one of the random masking or channel masking, and a threshold $\mu$ is used to control the probability of the two masking strategies being selected. Table 9 presents the experimental results, showing that the hybrid masking strategy achieves the highest accuracy and lowest standard deviation in both experiments. This suggests that the integration of different masking approaches significantly improves the model performance and stability. In addition, parallel masking increases the richness of augmented samples by combining different strategies, which is slightly better than using a single masking approach. However, this strategy makes the model training process unstable, resulting in large standard deviations. In summary, our proposed hybrid masking strategy is highly flexible and can be extended to various datasets by integrating multiple masking strategies using different ratios, which is exceptionally suitable for data with rich semantic information. This strategy effectively generates more diverse masked samples, encouraging the model to comprehensively capture the inherent relationships of the data.
Table 10. Subject-dependent and subject-independent EEG-based emotion recognition performance on SEED and SEED-IV.
Methods Subject-Dependent Subject-Independent
SEED SEED-IV SEED SEED-IV
SVM (Suykens and Vandewalle, 1999) 83.99 / 09.72 56.61 / 20.02 56.73 / 16.29 37.99 / 12.52
TCA (Pan et al., 2010) - - 63.38 / 14.88 37.01 / 10.47
GFK (Gong et al., 2012) - - 71.31 / 14.09 44.04 / 09.31
M3D (Luo et al., 2025) - - 84.57 / 09.49 60.94 / 08.84
Deep Learning
DBN (Zheng and Lu, 2015) 86.08 / 08.34 66.77 / 07.38 - -
DANN (Ganin et al., 2016) 83.99 / 09.72 56.73 / 16.29 56.61 / 20.02 37.99 / 12.52
DAN (Li et al., 2018a) - - 83.81 / 08.56 58.87 / 08.13
DGCNN (Song et al., 2018) 90.40 / 08.49 69.88 / 16.29 79.95 / 09.02 52.82 / 09.23
BiDANN (Li et al., 2018b) 92.38 / 07.04 70.29 / 12.63 84.14 / 06.87 65.59 / 10.39
BiHDM (Li et al., 2020) 93.12 / 06.06 74.35 / 14.09 85.40 / 07.53 69.03 / 08.66
RGNN (Zhong et al., 2020) 94.24 / 05.95 79.37 / 10.54 85.30 / 06.72 73.84 / 08.02
PR-PL (Zhou et al., 2023) 94.84 / 09.16 83.33 / 10.61 93.06 / 05.12 81.32 / 08.53
PGCN (Jin et al., 2024) 96.93 / 05.11 82.24 / 14.85 84.59 / 08.68 73.69 / 07.16
Self-Supervised Learning
SimCLR (Chen et al., 2020) 81.79 / 11.15 52.47 / 11.57 63.45 / 15.96 50.07 / 11.17
MoCo (He et al., 2020) 76.58 / 10.72 49.40 / 10.99 58.26 / 15.05 46.19 / 10.04
SSL-EEG (Xie et al., 2021) 83.32 / 09.20 63.59 / 19.82 67.52 / 12.73 53.62 / 08.47
GMSS (Li et al., 2022a) 89.18 / 09.74 65.61 / 17.33 76.04 / 11.91 62.13 / 08.33
SCMM (P-T) 92.19 / 07.68 81.41 / 08.17 83.84 / 06.22 75.28 / 06.37
SCMM (E-E) 93.02 / 06.67 81.60 / 08.12 84.88 / 05.85 76.09 / 06.77
Appendix E Within-Dataset Validation
To evaluate the model performance of SCMM for EEG-based emotion recognition within a single dataset, we conduct additional experiments on SEED and SEED-IV based on two experimental protocols: subject-dependent and subject-independent. In the experiments, we compare the proposed SCMM with three different types of models: machine learning, deep learning, and self-supervised learning. Additionally, we explore the model performance of SCMM using two different training strategies: pre-training-fine-tuning (P-T) and end-to-end (E-E). Table 10 presents the experimental results, indicating that our model achieves competitive results compared with various advanced deep learning models and significantly outperforms traditional machine learning methods and self-supervised learning models. In addition, the SCMM model using the end-to-end training strategy is better than the pre-training-fine-tuning strategy due to the introduction of the emotion classifier for joint learning. In summary, comprehensive experimental results demonstrate the effectiveness of the proposed SCMM in within-dataset EEG-based emotion recognition tasks.
Appendix F Full Results
Tables 11 to 14 present the full experimental results of SCMM.
Table 11. Full results on SEED, SEED-IV, and DEAP under same-class and different-class cross-corpus conditions.
Scenarios Accuracy Precision Recall F1 Score AUROC AUPRC
Same-Class
SEED-IV3 $\rightarrow$ SEED3 91.61 / 07.56 93.38 / 05.71 91.35 / 07.89 91.02 / 08.40 95.45 / 05.82 91.90 / 09.71
SEED3 $\rightarrow$ SEED-IV3 87.24 / 08.35 87.10 / 09.02 85.58 / 09.59 84.39 / 10.07 89.91 / 07.78 86.81 / 10.12
Different-Class
SEED-IV4 $\rightarrow$ SEED3 91.26 / 07.91 92.62 / 06.79 91.02 / 08.14 90.77 / 08.69 95.59 / 06.12 92.28 / 09.46
SEED3 $\rightarrow$ SEED-IV4 80.89 / 08.69 78.02 / 08.39 82.84 / 06.68 77.08 / 07.72 89.98 / 06.09 83.96 / 08.68
DEAP3 $\rightarrow$ SEED3 91.70 / 08.07 92.94 / 06.85 91.48 / 08.29 91.15 / 08.88 95.45 / 06.64 92.34 / 10.49
SEED3 $\rightarrow$ DEAP (Valence) 73.96 / 06.75 68.11 / 08.89 65.77 / 07.75 65.68 / 08.75 74.15 / 07.34 72.12 / 07.68
SEED3 $\rightarrow$ DEAP (Arousal) 72.66 / 05.67 70.97 / 05.91 68.90 / 07.11 68.84 / 07.57 75.77 / 07.33 74.79 / 07.50
Table 12. Full results of ablation study on SEED and SEED-IV under same-class and different-class cross-corpus conditions.
Scenarios Accuracy Precision Recall F1 Score AUROC AUPRC
w/o Soft Contrastive Loss $\mathcal{L}_{\rm C}$
SEED-IV3 $\rightarrow$ SEED3 89.68 / 09.32 91.81 / 07.76 89.39 / 09.58 89.20 / 09.97 94.71 / 05.95 91.73 / 09.32
SEED3 $\rightarrow$ SEED-IV3 84.24 / 11.90 84.79 / 14.99 83.84 / 10.79 81.94 / 14.11 90.65 / 10.10 85.69 / 15.93
SEED-IV4 $\rightarrow$ SEED3 89.45 / 09.10 91.23 / 08.34 89.21 / 09.32 89.13 / 09.58 95.11 / 05.86 91.24 / 10.77
SEED3 $\rightarrow$ SEED-IV4 77.24 / 09.14 73.21 / 14.36 73.55 / 12.89 70.24 / 14.29 88.57 / 09.82 80.79 / 16.07
w/o Reconstruction Loss $\mathcal{L}_{\rm R}$
SEED-IV3 $\rightarrow$ SEED3 90.73 / 08.48 93.22 / 06.17 90.46 / 08.76 90.26 / 09.07 94.57 / 06.58 90.52 / 10.06
SEED3 $\rightarrow$ SEED-IV3 85.07 / 11.05 85.26 / 14.01 81.59 / 13.58 80.88 / 15.46 88.35 / 15.99 85.16 / 18.10
SEED-IV4 $\rightarrow$ SEED3 90.96 / 08.36 92.43 / 07.24 90.70 / 08.57 90.41 / 09.07 94.36 / 06.39 89.98 / 10.40
SEED3 $\rightarrow$ SEED-IV4 78.32 / 07.19 73.07 / 14.36 74.59 / 12.07 72.01 / 13.92 87.02 / 12.49 82.09 / 16.17
Full Model
SEED-IV3 $\rightarrow$ SEED3 91.61 / 07.56 93.38 / 05.71 91.35 / 07.89 91.02 / 08.40 95.45 / 05.82 91.90 / 09.71
SEED3 $\rightarrow$ SEED-IV3 87.24 / 08.35 87.10 / 09.02 85.58 / 09.59 84.39 / 10.07 89.91 / 07.78 86.81 / 10.12
SEED-IV4 $\rightarrow$ SEED3 91.26 / 07.91 92.62 / 06.79 91.02 / 08.14 90.77 / 08.69 95.59 / 06.12 92.28 / 09.46
SEED3 $\rightarrow$ SEED-IV4 80.89 / 08.69 78.02 / 08.39 82.84 / 06.68 77.08 / 07.72 89.98 / 06.09 83.96 / 08.68
Table 13. Full results of model performance with limited labeled data for fine-tuning on SEED and SEED-IV under same-class and different-class cross-corpus conditions.
Scenarios Accuracy Precision Recall F1 Score AUROC AUPRC
1% of Labeled Data
SEED-IV3 $\rightarrow$ SEED3 74.98 / 17.00 74.52 / 21.19 74.69 / 17.11 70.64 / 21.53 93.76 / 06.20 89.32 / 09.68
SEED-IV4 $\rightarrow$ SEED3 73.20 / 16.09 73.41 / 21.28 72.93 / 16.23 69.14 / 20.07 94.32 / 04.69 90.31 / 07.53
SEED3 $\rightarrow$ SEED-IV3 71.35 / 16.29 72.22 / 18.40 69.37 / 15.47 66.23 / 18.19 80.18 / 14.09 74.07 / 17.64
SEED3 $\rightarrow$ SEED-IV4 60.59 / 23.72 55.23 / 29.74 58.65 / 22.49 52.48 / 27.81 82.19 / 13.44 70.54 / 19.49
5% of Labeled Data
SEED-IV3 $\rightarrow$ SEED3 95.23 / 06.98 96.20 / 04.72 95.10 / 07.14 94.86 / 07.88 99.67 / 00.38 99.45 / 00.62
SEED-IV4 $\rightarrow$ SEED3 90.51 / 12.43 92.90 / 08.38 90.27 / 12.70 89.04 / 15.29 98.23 / 02.25 96.76 / 04.34
SEED3 $\rightarrow$ SEED-IV3 87.20 / 16.12 88.31 / 14.88 86.08 / 15.69 85.35 / 17.64 91.93 / 10.70 89.14 / 13.71
SEED3 $\rightarrow$ SEED-IV4 86.01 / 09.76 84.86 / 14.77 82.69 / 11.35 81.29 / 13.75 95.14 / 03.37 92.56 / 06.89
10% of Labeled Data
SEED-IV3 $\rightarrow$ SEED3 99.86 / 00.31 99.87 / 00.30 99.86 / 00.31 99.86 / 00.31 100.00 / 00.00 100.00 / 00.00
SEED-IV4 $\rightarrow$ SEED3 99.72 / 00.80 99.74 / 00.74 99.71 / 00.81 99.72 / 00.80 99.87 / 00.46 99.78 / 00.77
SEED3 $\rightarrow$ SEED-IV3 90.04 / 16.94 91.00 / 17.05 90.09 / 15.87 89.32 / 18.21 94.29 / 10.86 92.55 / 13.98
SEED3 $\rightarrow$ SEED-IV4 94.43 / 05.60 95.66 / 04.71 92.18 / 08.23 92.19 / 08.82 98.12 / 02.13 96.03 / 05.45
20% of Labeled Data
SEED-IV3 $\rightarrow$ SEED3 100.00 / 00.00 100.00 / 00.00 100.00 / 00.00 100.00 / 00.00 100.00 / 00.00 100.00 / 00.00
SEED-IV4 $\rightarrow$ SEED3 100.00 / 00.00 100.00 / 00.00 100.00 / 00.00 100.00 / 00.00 100.00 / 00.00 100.00 / 00.00
SEED3 $\rightarrow$ SEED-IV3 94.16 / 08.52 95.20 / 06.36 93.66 / 08.00 93.72 / 08.35 97.59 / 03.38 96.53 / 04.85
SEED3 $\rightarrow$ SEED-IV4 98.91 / 02.62 98.82 / 02.97 98.41 / 03.73 98.52 / 03.65 98.97 / 03.20 98.13 / 05.64
Table 14. Full results of hyperparameter sensitivity analysis on SEED and SEED-IV under the same-class cross-corpus condition.
Hyperparameters SEED-IV3 $\rightarrow$ SEED3 SEED3 $\rightarrow$ SEED-IV3
Masking Ratio $r$
@@ Appendix F Full Results
Threshold $\mu$
@@
Metric Function $Dist(\cdot,\cdot)$
@@
Upper Bound $\alpha$
0.25 91.21 / 08.54 -
0.5 91.61 / 07.56 -
0.75 91.12 / 08.25 -
1 90.50 / 08.00 -
Sharpness $\tau_{\rm s}$
0.01 90.93 / 09.13 -
0.05 91.61 / 07.56 -
0.1 91.08 / 09.00 -
0.2 90.59 / 07.40 -
0.5 90.63 / 09.16 -
1 90.09 / 08.62 -
2 89.58 / 09.26 -
Temperature $\tau_{\rm c}$
@@
Pre-training Batch Size
8 91.41 / 08.42 86.64 / 11.77
16 91.29 / 08.90 86.33 / 12.04
32 91.18 / 08.23 86.77 / 12.72
64 90.35 / 08.37 86.85 / 10.96
128 91.00 / 08.22 85.74 / 13.42
256 91.61 / 07.56 87.24 / 08.35
512 90.04 / 08.04 86.74 / 11.14